# Hui Xiong2, Bin Yang1

### v1

@@ Appendix C Full Experimental Results
Dataset Metric CATCH Modern iTrans DualTF ATrans DC TsNet Patch DLin NLin AE Ocsvm IF PCA HBOS
CICIDS A-P 0.002 0.001 0.002 0.002 0.001 0.002 0.002 0.001 0.002 0.001 0.001 0.001 0.003 0.001 0.003
R-A-P 0.002 0.002 0.002 0.001 0.001 0.001 0.002 0.002 0.002 0.002 0.001 0.364 0.002 0.001 0.002
V-PR 0.003 0.002 0.002 0.002 0.001 0.002 0.002 0.002 0.002 0.002 0.002 0.389 0.003 0.001 0.002
CalIt2 A-P 0.114 0.054 0.106 0.057 0.045 0.035 0.078 0.116 0.097 0.054 0.084 0.095 0.080 0.073 0.080
R-A-P 0.124 0.070 0.111 0.088 0.089 0.087 0.092 0.115 0.097 0.073 0.096 0.109 0.095 0.106 0.113
V-PR 0.121 0.070 0.110 0.082 0.083 0.083 0.090 0.115 0.095 0.072 0.097 0.109 0.091 0.103 0.108
Credit A-P 0.101 0.088 0.042 0.023 0.007 0.002 0.091 0.089 0.081 0.087 0.040 0.053 0.074 0.029 0.173
R-A-P 0.053 0.054 0.024 0.009 0.002 0.002 0.056 0.054 0.053 0.053 0.020 0.038 0.040 0.009 0.081
V-PR 0.051 0.051 0.024 0.011 0.002 0.002 0.053 0.051 0.050 0.051 0.020 0.037 0.039 0.012 0.076
GECCO A-P 0.418 0.447 0.096 0.130 0.013 0.012 0.410 0.400 0.349 0.304 0.206 0.039 0.052 0.234 0.199
R-A-P 0.473 0.459 0.134 0.050 0.022 0.020 0.428 0.444 0.416 0.372 0.033 0.098 0.085 0.047 0.032
V-PR 0.465 0.461 0.128 0.049 0.022 0.019 0.429 0.439 0.406 0.363 0.033 0.101 0.083 0.046 0.033
Genesis A-P 0.249 0.015 0.019 0.051 0.058 0.010 0.036 0.013 0.017 0.011 0.055 0.007 0.005 0.011 0.059
R-A-P 0.384 0.015 0.021 0.101 0.101 0.012 0.038 0.013 0.016 0.013 0.047 0.506 0.011 0.015 0.090
V-PR 0.371 0.015 0.020 0.097 0.095 0.012 0.037 0.013 0.016 0.014 0.047 0.506 0.010 0.015 0.087
MSL A-P 0.167 0.146 0.151 0.156 0.107 0.107 0.146 0.157 0.147 0.140 0.148 0.153 0.114 0.157 0.132
R-A-P 0.260 0.224 0.227 0.220 0.165 0.157 0.231 0.242 0.225 0.222 0.200 0.185 0.174 0.203 0.190
V-PR 0.256 0.220 0.224 0.218 0.162 0.156 0.227 0.237 0.221 0.218 0.199 0.185 0.173 0.200 0.189
NYC A-P 0.076 0.020 0.033 0.040 0.045 0.034 0.060 0.046 0.046 0.062 0.025 0.020 0.022 0.046 0.019
R-A-P 0.120 0.038 0.051 0.065 0.096 0.045 0.069 0.059 0.061 0.066 0.046 0.333 0.047 0.076 0.031
V-PR 0.114 0.037 0.051 0.063 0.087 0.047 0.070 0.059 0.061 0.065 0.045 0.318 0.045 0.072 0.032
PSM A-P 0.434 0.385 0.383 0.411 0.298 0.281 0.391 0.378 0.371 0.376 0.465 0.418 0.334 0.468 0.394
R-A-P 0.435 0.383 0.386 0.353 0.293 0.283 0.395 0.379 0.372 0.378 0.420 0.369 0.334 0.423 0.369
V-PR 0.436 0.384 0.387 0.354 0.293 0.283 0.395 0.380 0.373 0.379 0.420 0.370 0.334 0.423 0.370
SMD A-P 0.172 0.130 0.146 0.069 0.046 0.043 0.141 0.147 0.139 0.141 0.188 0.104 0.122 0.128 0.145
R-A-P 0.159 0.130 0.145 0.070 0.055 0.046 0.140 0.152 0.145 0.145 0.182 0.080 0.099 0.109 0.088
V-PR 0.159 0.130 0.145 0.070 0.054 0.046 0.140 0.152 0.144 0.145 0.181 0.081 0.099 0.109 0.088
Contextual4.9 A-P 0.754 0.090 0.754 0.086 0.070 0.054 0.773 0.640 0.274 0.055 0.732 0.522 0.323 0.066 0.044
R-A-P 0.585 0.060 0.585 0.075 0.055 0.055 0.614 0.432 0.135 0.054 0.556 0.293 0.225 0.054 0.047
V-PR 0.562 0.074 0.559 0.094 0.068 0.067 0.586 0.423 0.148 0.066 0.534 0.292 0.230 0.068 0.057
Contextual7.2 A-P 0.770 0.106 0.756 0.087 0.099 0.074 0.767 0.672 0.309 0.090 0.721 0.542 0.374 0.105 0.075
R-A-P 0.612 0.083 0.594 0.078 0.078 0.079 0.608 0.477 0.166 0.075 0.546 0.324 0.260 0.079 0.067
V-PR 0.594 0.100 0.579 0.097 0.094 0.094 0.590 0.469 0.183 0.090 0.533 0.333 0.271 0.098 0.081
Global4.8 A-P 0.978 0.552 0.939 0.087 0.064 0.051 0.959 0.959 0.927 0.099 0.971 0.981 0.435 0.136 0.113
R-A-P 0.957 0.363 0.881 0.068 0.055 0.050 0.918 0.918 0.858 0.077 0.942 0.961 0.404 0.110 0.063
V-PR 0.901 0.361 0.834 0.092 0.069 0.062 0.866 0.864 0.808 0.089 0.884 0.903 0.394 0.140 0.074
Global7.2 A-P 0.974 0.517 0.921 0.088 0.098 0.074 0.947 0.960 0.912 0.134 0.973 0.977 0.492 0.182 0.106
R-A-P 0.949 0.346 0.854 0.082 0.084 0.074 0.902 0.923 0.834 0.106 0.948 0.953 0.429 0.145 0.090
V-PR 0.905 0.347 0.818 0.099 0.101 0.089 0.860 0.876 0.796 0.122 0.899 0.904 0.426 0.182 0.107
Seasonal4.8 A-P 0.975 0.055 0.865 0.476 0.158 0.128 0.889 0.776 0.622 0.063 0.884 0.732 0.727 0.061 0.058
R-A-P 0.994 0.087 0.739 0.719 0.162 0.223 0.750 0.681 0.540 0.099 0.718 0.604 0.679 0.092 0.088
V-PR 0.987 0.084 0.755 0.676 0.165 0.210 0.763 0.695 0.551 0.099 0.737 0.621 0.679 0.091 0.085
Seasonal7.7 A-P 0.963 0.092 0.881 0.083 0.164 0.148 0.906 0.773 0.645 0.109 0.900 0.734 0.733 0.116 0.080
R-A-P 0.978 0.148 0.808 0.139 0.192 0.260 0.805 0.723 0.599 0.163 0.758 0.645 0.752 0.169 0.132
V-PR 0.963 0.141 0.819 0.134 0.194 0.246 0.823 0.732 0.605 0.159 0.782 0.658 0.743 0.160 0.132
Shapelet4.9 A-P 0.774 0.065 0.633 0.124 0.066 0.142 0.683 0.499 0.382 0.053 0.659 0.482 0.282 0.094 0.043
R-A-P 0.940 0.089 0.564 0.199 0.092 0.153 0.588 0.463 0.314 0.090 0.556 0.410 0.319 0.138 0.070
V-PR 0.887 0.089 0.571 0.195 0.090 0.147 0.599 0.465 0.322 0.089 0.568 0.422 0.311 0.135 0.071
Shapelet7.4 A-P 0.774 0.087 0.651 0.089 0.109 0.133 0.708 0.542 0.449 0.089 0.715 0.537 0.295 0.136 0.062
R-A-P 0.921 0.132 0.608 0.149 0.143 0.169 0.639 0.528 0.416 0.137 0.619 0.488 0.352 0.241 0.104
V-PR 0.877 0.133 0.616 0.148 0.137 0.164 0.653 0.534 0.422 0.136 0.634 0.496 0.343 0.224 0.101
Mixture5.7 A-P 0.368 0.196 0.238 0.060 0.060 0.088 0.268 0.273 0.172 0.104 0.068 0.050 0.591 0.082 0.051
R-A-P 0.509 0.276 0.349 0.110 0.092 0.146 0.363 0.369 0.204 0.160 0.091 0.125 0.579 0.092 0.122
V-PR 0.477 0.264 0.339 0.106 0.090 0.143 0.352 0.356 0.197 0.157 0.089 0.122 0.577 0.090 0.125
Mixture5.9 A-P 0.355 0.182 0.311 0.091 0.091 0.133 0.365 0.298 0.197 0.167 0.128 0.090 0.593 0.090 0.089
R-A-P 0.490 0.253 0.386 0.160 0.152 0.199 0.403 0.364 0.239 0.243 0.165 0.250 0.627 0.154 0.226
V-PR 0.468 0.247 0.375 0.157 0.150 0.192 0.393 0.357 0.233 0.236 0.162 0.246 0.619 0.151 0.221
Trend4.8 A-P 0.377 0.215 0.254 0.062 0.048 0.063 0.274 0.228 0.155 0.133 0.056 0.044 0.615 0.054 0.054
R-A-P 0.538 0.343 0.368 0.093 0.092 0.093 0.400 0.385 0.241 0.223 0.090 0.330 0.589 0.092 0.239
V-PR 0.516 0.329 0.360 0.092 0.097 0.092 0.389 0.375 0.237 0.216 0.089 0.318 0.588 0.091 0.233
Trend7.8 A-P 0.510 0.520 0.327 0.126 0.106 0.078 0.384 0.501 0.459 0.359 0.091 0.075 0.511 0.075 0.082
R-A-P 0.665 0.514 0.399 0.173 0.171 0.128 0.435 0.521 0.429 0.391 0.154 0.261 0.550 0.135 0.250
V-PR 0.637 0.515 0.386 0.169 0.166 0.125 0.424 0.512 0.429 0.389 0.152 0.252 0.543 0.133 0.240
$1^{st}$ Count 29 2 0 0 0 0 3 1 0 0 3 11 10 1 3
Table 7: Average Acc (Accuracy), P (Precision), R (Recall) and F1 (F1-score) accuracy measures for all datasets. The best results are highlighted in bold, and the second-best results are underlined.
Dataset Metric CATCH Modern iTrans DualTF ATrans DC TsNet Patch DLin NLin AE Ocsvm IF PCA HBOS
CICIDS Acc 0.758 0.750 0.762 0.565 0.745 0.750 0.750 0.751 0.757 0.754 0.994 0.082 0.852 0.656 0.913
P 0.003 0.001 0.002 0.001 0.001 0.001 0.001 0.001 0.002 0.001 0.000 0.001 0.000 0.001 0.000
R 0.759 0.278 0.481 0.595 0.253 0.266 0.354 0.380 0.506 0.380 0.000 0.987 0.063 0.405 0.013
F1 0.006 0.002 0.004 0.003 0.002 0.002 0.003 0.003 0.004 0.003 0.000 0.002 0.001 0.002 0.000
CalIt2 Acc 0.946 0.889 0.930 0.910 0.944 0.841 0.946 0.889 0.887 0.844 0.960 0.867 0.968 0.884 0.886
P 0.138 0.074 0.124 0.073 0.085 0.034 0.104 0.091 0.083 0.066 0.067 0.089 0.125 0.084 0.075
R 0.162 0.243 0.230 0.176 0.095 0.162 0.108 0.311 0.284 0.324 0.027 0.378 0.014 0.297 0.257
F1 0.149 0.114 0.161 0.103 0.090 0.056 0.106 0.141 0.128 0.109 0.038 0.144 0.024 0.131 0.117
Credit Acc 0.981 0.981 0.951 0.688 0.847 0.716 0.981 0.980 0.981 0.980 0.994 0.893 0.996 0.868 0.847
P 0.059 0.058 0.026 0.003 0.001 0.001 0.058 0.058 0.057 0.057 0.047 0.013 0.156 0.009 0.009
R 0.758 0.740 0.843 0.596 0.139 0.238 0.740 0.749 0.726 0.744 0.135 0.901 0.283 0.731 0.915
F1 0.110 0.107 0.051 0.006 0.003 0.003 0.107 0.107 0.105 0.107 0.069 0.026 0.201 0.017 0.018
GECCO Acc 0.984 0.984 0.981 0.612 0.983 0.979 0.984 0.989 0.981 0.987 0.888 0.055 0.989 0.240 0.590
P 0.380 0.373 0.173 0.009 0.012 0.008 0.379 0.475 0.332 0.384 0.021 0.011 0.214 0.014 0.014
R 0.818 0.779 0.207 0.340 0.008 0.008 0.804 0.585 0.781 0.460 0.215 0.993 0.012 0.988 0.542
F1 0.518 0.504 0.189 0.018 0.010 0.008 0.516 0.524 0.466 0.418 0.039 0.022 0.023 0.027 0.027
Genesis Acc 0.992 0.965 0.986 0.970 0.931 0.991 0.991 0.987 0.987 0.972 0.987 0.424 0.983 0.888 0.771
P 0.116 0.015 0.065 0.066 0.053 0.016 0.119 0.075 0.055 0.013 0.047 0.007 0.006 0.017 0.017
R 0.160 0.120 0.180 0.500 0.960 0.020 0.200 0.200 0.140 0.080 0.120 1.000 0.020 0.480 0.980
F1 0.134 0.027 0.095 0.116 0.100 0.018 0.149 0.109 0.079 0.022 0.068 0.014 0.009 0.033 0.033
MSL Acc 0.853 0.857 0.854 0.891 0.891 0.892 0.855 0.873 0.858 0.872 0.891 0.818 0.890 0.812 0.816
P 0.185 0.166 0.158 0.248 0.143 0.181 0.166 0.194 0.173 0.193 0.219 0.128 0.180 0.130 0.127
R 0.117 0.090 0.089 0.019 0.008 0.008 0.093 0.066 0.093 0.067 0.014 0.125 0.014 0.137 0.128
F1 0.143 0.117 0.114 0.035 0.015 0.016 0.119 0.098 0.121 0.100 0.026 0.126 0.025 0.133 0.127
NYC Acc 0.978 0.972 0.797 0.976 0.977 0.977 0.977 0.976 0.976 0.977 0.974 0.029 0.974 0.904 0.895
P 1.000 0.000 0.034 0.111 0.333 0.250 0.000 0.000 0.000 0.000 0.059 0.022 0.056 0.059 0.000
R 0.010 0.000 0.293 0.010 0.010 0.010 0.000 0.000 0.000 0.000 0.010 0.980 0.010 0.222 0.000
F1 0.020 0.000 0.061 0.019 0.020 0.019 0.000 0.000 0.000 0.000 0.017 0.043 0.017 0.094 0.000
PSM Acc 0.730 0.729 0.728 0.697 0.719 0.615 0.726 0.729 0.726 0.728 0.715 0.737 0.723 0.682 0.736
P 0.624 0.653 0.600 0.461 0.396 0.278 0.591 0.653 0.658 0.644 0.444 0.548 0.802 0.427 0.544
R 0.064 0.047 0.058 0.559 0.025 0.243 0.048 0.047 0.025 0.047 0.110 0.299 0.004 0.421 0.302
F1 0.116 0.089 0.105 0.506 0.046 0.259 0.088 0.088 0.047 0.087 0.176 0.387 0.007 0.424 0.389
SMD Acc 0.918 0.931 0.914 0.863 0.952 0.774 0.931 0.934 0.943 0.943 0.959 0.865 0.958 0.860 0.841
P 0.194 0.151 0.162 0.091 0.112 0.042 0.176 0.185 0.197 0.201 0.706 0.107 0.454 0.117 0.095
R 0.305 0.145 0.255 0.257 0.022 0.201 0.181 0.173 0.124 0.126 0.007 0.304 0.020 0.359 0.332
F1 0.237 0.148 0.198 0.135 0.037 0.069 0.178 0.179 0.152 0.155 0.014 0.158 0.039 0.176 0.148
Contextual4.9 Acc 0.961 0.734 0.957 0.713 0.641 0.618 0.734 0.888 0.943 0.742 0.965 0.888 0.956 0.848 0.865
P 0.584 0.067 0.555 0.044 0.051 0.049 0.101 0.264 0.446 0.119 0.670 0.234 0.577 0.072 0.029
R 0.733 0.340 0.696 0.235 0.356 0.364 0.559 0.713 0.656 0.656 0.591 0.559 0.393 0.174 0.053
F1 0.650 0.112 0.618 0.075 0.089 0.086 0.172 0.385 0.531 0.201 0.628 0.330 0.467 0.102 0.037
Contextual7.2 Acc 0.960 0.721 0.937 0.603 0.702 0.611 0.739 0.886 0.893 0.729 0.956 0.877 0.939 0.818 0.850
P 0.744 0.095 0.548 0.085 0.080 0.067 0.157 0.352 0.373 0.160 0.775 0.306 0.633 0.103 0.053
R 0.686 0.336 0.719 0.461 0.297 0.342 0.600 0.700 0.717 0.650 0.556 0.556 0.350 0.197 0.064
F1 0.714 0.148 0.622 0.143 0.126 0.112 0.249 0.468 0.491 0.257 0.647 0.394 0.451 0.135 0.058
Global4.8 Acc 0.993 0.896 0.994 0.833 0.174 0.639 0.989 0.995 0.993 0.895 0.980 0.903 0.977 0.698 0.872
P 0.936 0.277 1.000 0.071 0.048 0.046 0.900 1.000 0.977 0.290 0.747 0.331 0.766 0.109 0.086
R 0.921 0.721 0.875 0.204 0.854 0.329 0.858 0.888 0.871 0.825 0.875 0.996 0.750 0.738 0.175
F1 0.929 0.400 0.933 0.105 0.090 0.080 0.878 0.940 0.921 0.429 0.806 0.496 0.758 0.190 0.116
Global7.2 Acc 0.978 0.849 0.972 0.326 0.232 0.665 0.987 0.984 0.985 0.900 0.978 0.915 0.970 0.644 0.843
[…]

### v3

@@ Appendix C Full Experimental Results
Dataset Metric CATCH Modern iTrans DualTF ATrans DC TsNet Patch DLin NLin AE TFAD Ocsvm IF PCA HBOS
CICIDS A-P 0.002 0.001 0.002 0.002 0.001 0.002 0.002 0.001 0.002 0.001 0.001 0.003 0.001 0.003 0.001 0.003
R-A-P 0.002 0.002 0.002 0.001 0.001 0.001 0.002 0.002 0.002 0.002 0.001 0.002 0.364 0.002 0.001 0.002
V-PR 0.003 0.002 0.002 0.002 0.001 0.002 0.002 0.002 0.002 0.002 0.002 0.002 0.389 0.003 0.001 0.002
CalIt2 A-P 0.114 0.054 0.106 0.057 0.045 0.035 0.078 0.116 0.097 0.054 0.084 0.030 0.095 0.080 0.073 0.080
R-A-P 0.124 0.070 0.111 0.088 0.089 0.087 0.092 0.115 0.097 0.073 0.096 0.060 0.109 0.095 0.106 0.113
V-PR 0.121 0.070 0.110 0.082 0.083 0.083 0.090 0.115 0.095 0.072 0.097 0.051 0.109 0.091 0.103 0.108
Credit A-P 0.101 0.088 0.042 0.023 0.007 0.002 0.091 0.089 0.081 0.087 0.040 0.002 0.053 0.074 0.029 0.173
R-A-P 0.053 0.054 0.024 0.009 0.002 0.002 0.056 0.054 0.053 0.053 0.020 0.002 0.038 0.040 0.009 0.081
V-PR 0.051 0.051 0.024 0.011 0.002 0.002 0.053 0.051 0.050 0.051 0.020 0.002 0.037 0.039 0.012 0.076
GECCO A-P 0.418 0.447 0.096 0.130 0.013 0.012 0.410 0.400 0.349 0.304 0.206 0.012 0.039 0.052 0.234 0.199
R-A-P 0.473 0.459 0.134 0.050 0.022 0.020 0.428 0.444 0.416 0.372 0.033 0.019 0.098 0.085 0.047 0.032
V-PR 0.465 0.461 0.128 0.049 0.022 0.019 0.429 0.439 0.406 0.363 0.033 0.018 0.101 0.083 0.046 0.033
Genesis A-P 0.249 0.015 0.019 0.051 0.058 0.010 0.036 0.013 0.017 0.011 0.055 0.004 0.007 0.005 0.011 0.059
R-A-P 0.384 0.015 0.021 0.101 0.101 0.012 0.038 0.013 0.016 0.013 0.047 0.003 0.506 0.011 0.015 0.090
V-PR 0.371 0.015 0.020 0.097 0.095 0.012 0.037 0.013 0.016 0.014 0.047 0.003 0.506 0.010 0.015 0.087
MSL A-P 0.167 0.146 0.151 0.156 0.107 0.107 0.146 0.157 0.147 0.140 0.148 0.106 0.153 0.114 0.157 0.132
R-A-P 0.260 0.224 0.227 0.220 0.165 0.157 0.231 0.242 0.225 0.222 0.200 0.162 0.185 0.174 0.203 0.190
V-PR 0.256 0.220 0.224 0.218 0.162 0.156 0.227 0.237 0.221 0.218 0.199 0.159 0.185 0.173 0.200 0.189
NYC A-P 0.076 0.020 0.033 0.040 0.045 0.034 0.060 0.046 0.046 0.062 0.025 0.023 0.020 0.022 0.046 0.019
R-A-P 0.120 0.038 0.051 0.065 0.096 0.045 0.069 0.059 0.061 0.066 0.046 0.046 0.333 0.047 0.076 0.031
V-PR 0.114 0.037 0.051 0.063 0.087 0.047 0.070 0.059 0.061 0.065 0.045 0.047 0.318 0.045 0.072 0.032
PSM A-P 0.434 0.385 0.383 0.411 0.298 0.281 0.391 0.378 0.371 0.376 0.465 0.279 0.418 0.334 0.468 0.394
R-A-P 0.435 0.383 0.386 0.353 0.293 0.283 0.395 0.379 0.372 0.378 0.420 0.291 0.369 0.334 0.423 0.369
V-PR 0.436 0.384 0.387 0.354 0.293 0.283 0.395 0.380 0.373 0.379 0.420 0.291 0.370 0.334 0.423 0.370
SMD A-P 0.172 0.130 0.146 0.069 0.046 0.043 0.141 0.147 0.139 0.141 0.188 0.042 0.104 0.122 0.128 0.145
R-A-P 0.159 0.130 0.145 0.070 0.055 0.046 0.140 0.152 0.145 0.145 0.182 0.052 0.080 0.099 0.109 0.088
V-PR 0.159 0.130 0.145 0.070 0.054 0.046 0.140 0.152 0.144 0.145 0.181 0.052 0.081 0.099 0.109 0.088
ASD A-P 0.231 0.158 0.164 0.095 0.052 0.049 0.245 0.174 0.170 0.132 0.203 0.052 0.142 0.195 0.144 0.133
R-A-P 0.269 0.170 0.197 0.103 0.082 0.090 0.254 0.202 0.191 0.139 0.208 0.089 0.102 0.197 0.154 0.132
V-PR 0.261 0.170 0.194 0.103 0.080 0.090 0.253 0.200 0.188 0.139 0.208 0.091 0.103 0.197 0.149 0.130
Contextual4.9 A-P 0.754 0.090 0.754 0.086 0.070 0.054 0.773 0.640 0.274 0.055 0.732 0.049 0.522 0.323 0.066 0.044
R-A-P 0.585 0.060 0.585 0.075 0.055 0.055 0.614 0.432 0.135 0.054 0.556 0.025 0.293 0.225 0.054 0.047
V-PR 0.562 0.074 0.559 0.094 0.068 0.067 0.586 0.423 0.148 0.066 0.534 0.047 0.292 0.230 0.068 0.057
Contextual7.2 A-P 0.770 0.106 0.756 0.087 0.099 0.074 0.767 0.672 0.309 0.090 0.721 0.109 0.542 0.374 0.105 0.075
R-A-P 0.612 0.083 0.594 0.078 0.078 0.079 0.608 0.477 0.166 0.075 0.546 0.082 0.324 0.260 0.079 0.067
V-PR 0.594 0.100 0.579 0.097 0.094 0.094 0.590 0.469 0.183 0.090 0.533 0.097 0.333 0.271 0.098 0.081
Global4.8 A-P 0.978 0.552 0.939 0.087 0.064 0.051 0.959 0.959 0.927 0.099 0.971 0.098 0.981 0.435 0.136 0.113
R-A-P 0.957 0.363 0.881 0.068 0.055 0.050 0.918 0.918 0.858 0.077 0.942 0.051 0.961 0.404 0.110 0.063
V-PR 0.901 0.361 0.834 0.092 0.069 0.062 0.866 0.864 0.808 0.089 0.884 0.062 0.903 0.394 0.140 0.074
Global7.2 A-P 0.974 0.517 0.921 0.088 0.098 0.074 0.947 0.960 0.912 0.134 0.973 0.074 0.977 0.492 0.182 0.106
R-A-P 0.949 0.346 0.854 0.082 0.084 0.074 0.902 0.923 0.834 0.106 0.948 0.060 0.953 0.429 0.145 0.090
V-PR 0.905 0.347 0.818 0.099 0.101 0.089 0.860 0.876 0.796 0.122 0.899 0.067 0.904 0.426 0.182 0.107
Seasonal4.8 A-P 0.975 0.055 0.865 0.476 0.158 0.128 0.889 0.776 0.622 0.063 0.884 0.084 0.732 0.727 0.061 0.058
R-A-P 0.994 0.087 0.739 0.719 0.162 0.223 0.750 0.681 0.540 0.099 0.718 0.152 0.604 0.679 0.092 0.088
V-PR 0.987 0.084 0.755 0.676 0.165 0.210 0.763 0.695 0.551 0.099 0.737 0.138 0.621 0.679 0.091 0.085
Seasonal7.7 A-P 0.963 0.092 0.881 0.083 0.164 0.148 0.906 0.773 0.645 0.109 0.900 0.133 0.734 0.733 0.116 0.080
R-A-P 0.978 0.148 0.808 0.139 0.192 0.260 0.805 0.723 0.599 0.163 0.758 0.195 0.645 0.752 0.169 0.132
V-PR 0.963 0.141 0.819 0.134 0.194 0.246 0.823 0.732 0.605 0.159 0.782 0.185 0.658 0.743 0.160 0.132
Shapelet4.9 A-P 0.774 0.065 0.633 0.124 0.066 0.142 0.683 0.499 0.382 0.053 0.659 0.102 0.482 0.282 0.094 0.043
R-A-P 0.940 0.089 0.564 0.199 0.092 0.153 0.588 0.463 0.314 0.090 0.556 0.144 0.410 0.319 0.138 0.070
V-PR 0.887 0.089 0.571 0.195 0.090 0.147 0.599 0.465 0.322 0.089 0.568 0.137 0.422 0.311 0.135 0.071
Shapelet7.4 A-P 0.774 0.087 0.651 0.089 0.109 0.133 0.708 0.542 0.449 0.089 0.715 0.118 0.537 0.295 0.136 0.062
R-A-P 0.921 0.132 0.608 0.149 0.143 0.169 0.639 0.528 0.416 0.137 0.619 0.175 0.488 0.352 0.241 0.104
V-PR 0.877 0.133 0.616 0.148 0.137 0.164 0.653 0.534 0.422 0.136 0.634 0.171 0.496 0.343 0.224 0.101
Mixture5.7 A-P 0.368 0.196 0.238 0.060 0.060 0.088 0.268 0.273 0.172 0.104 0.068 0.059 0.050 0.591 0.082 0.051
R-A-P 0.509 0.276 0.349 0.110 0.092 0.146 0.363 0.369 0.204 0.160 0.091 0.163 0.125 0.579 0.092 0.122
V-PR 0.477 0.264 0.339 0.106 0.090 0.143 0.352 0.356 0.197 0.157 0.089 0.154 0.122 0.577 0.090 0.125
Mixture8.9 A-P 0.355 0.182 0.311 0.091 0.091 0.133 0.365 0.298 0.197 0.167 0.128 0.091 0.090 0.593 0.090 0.089
R-A-P 0.490 0.253 0.386 0.160 0.152 0.199 0.403 0.364 0.239 0.243 0.165 0.145 0.250 0.627 0.154 0.226
V-PR 0.468 0.247 0.375 0.157 0.150 0.192 0.393 0.357 0.233 0.236 0.162 0.149 0.246 0.619 0.151 0.221
Trend4.8 A-P 0.377 0.215 0.254 0.062 0.048 0.063 0.274 0.228 0.155 0.133 0.056 0.058 0.044 0.615 0.054 0.054
R-A-P 0.538 0.343 0.368 0.093 0.092 0.093 0.400 0.385 0.241 0.223 0.090 0.091 0.330 0.589 0.092 0.239
V-PR 0.516 0.329 0.360 0.092 0.097 0.092 0.389 0.375 0.237 0.216 0.089 0.091 0.318 0.588 0.091 0.233
Trend7.8 A-P 0.510 0.520 0.327 0.126 0.106 0.078 0.384 0.501 0.459 0.359 0.091 0.088 0.075 0.511 0.075 0.082
R-A-P 0.665 0.514 0.399 0.173 0.171 0.128 0.435 0.521 0.429 0.391 0.154 0.157 0.261 0.550 0.135 0.250
V-PR 0.637 0.515 0.386 0.169 0.166 0.125 0.424 0.512 0.429 0.389 0.152 0.150 0.252 0.543 0.133 0.240
$1^{st}$ Count 31 2 0 0 0 0 4 1 0 1 3 1 13 9 1 3
Table 8: Average P (Precision), R (Recall) and F1 (F1-score) accuracy measures for all datasets. The best results are highlighted in bold, and the second-best results are underlined.
Dataset Metric CATCH Modern iTrans DualTF ATrans DC TsNet Patch DLin NLin AE TFAD Ocsvm IF PCA HBOS
CICIDS P 0.003 0.001 0.002 0.001 0.001 0.001 0.001 0.001 0.002 0.001 0.000 0.001 0.001 0.000 0.001 0.000
R 0.759 0.278 0.481 0.595 0.253 0.266 0.354 0.380 0.506 0.380 0.000 0.038 0.987 0.063 0.405 0.013
F1 0.006 0.002 0.004 0.003 0.002 0.002 0.003 0.003 0.004 0.003 0.000 0.001 0.002 0.001 0.002 0.000
CalIt2 P 0.138 0.074 0.124 0.073 0.085 0.034 0.104 0.091 0.083 0.066 0.067 0.038 0.089 0.125 0.084 0.075
R 0.162 0.243 0.230 0.176 0.095 0.162 0.108 0.311 0.284 0.324 0.027 0.041 0.378 0.014 0.297 0.257
F1 0.149 0.114 0.161 0.103 0.090 0.056 0.106 0.141 0.128 0.109 0.038 0.039 0.144 0.024 0.131 0.117
Credit P 0.059 0.058 0.026 0.003 0.001 0.001 0.058 0.058 0.057 0.057 0.047 0.005 0.013 0.156 0.009 0.009
R 0.758 0.740 0.843 0.596 0.139 0.238 0.740 0.749 0.726 0.744 0.135 0.018 0.901 0.283 0.731 0.915
F1 0.110 0.107 0.051 0.006 0.003 0.003 0.107 0.107 0.105 0.107 0.069 0.007 0.026 0.201 0.017 0.018
GECCO P 0.380 0.373 0.173 0.009 0.012 0.008 0.379 0.475 0.332 0.384 0.021 0.009 0.011 0.214 0.014 0.014
R 0.818 0.779 0.207 0.340 0.008 0.008 0.804 0.585 0.781 0.460 0.215 0.004 0.993 0.012 0.988 0.542
F1 0.518 0.504 0.189 0.018 0.010 0.008 0.516 0.524 0.466 0.418 0.039 0.006 0.022 0.023 0.027 0.027
Genesis P 0.116 0.015 0.065 0.066 0.053 0.016 0.119 0.075 0.055 0.013 0.047 0.000 0.007 0.006 0.017 0.017
R 0.160 0.120 0.180 0.500 0.960 0.020 0.200 0.200 0.140 0.080 0.120 0.000 1.000 0.020 0.480 0.980
F1 0.134 0.027 0.095 0.116 0.100 0.018 0.149 0.109 0.079 0.022 0.068 0.000 0.014 0.009 0.033 0.033
MSL P 0.185 0.166 0.158 0.248 0.143 0.181 0.166 0.194 0.173 0.193 0.219 0.107 0.128 0.180 0.130 0.127
R 0.117 0.090 0.089 0.019 0.008 0.008 0.093 0.066 0.093 0.067 0.014 0.005 0.125 0.014 0.137 0.128
F1 0.143 0.117 0.114 0.035 0.015 0.016 0.119 0.098 0.121 0.100 0.026 0.010 0.126 0.025 0.133 0.127
NYC P 1.000 0.000 0.034 0.111 0.333 0.250 0.000 0.000 0.000 0.000 0.059 0.000 0.022 0.056 0.059 0.000
R 0.010 0.000 0.293 0.010 0.010 0.010 0.000 0.000 0.000 0.000 0.010 0.000 0.980 0.010 0.222 0.000
F1 0.020 0.000 0.061 0.019 0.020 0.019 0.000 0.000 0.000 0.000 0.017 0.000 0.043 0.017 0.094 0.000
PSM P 0.624 0.653 0.600 0.461 0.396 0.278 0.591 0.653 0.658 0.644 0.444 0.282 0.548 0.802 0.427 0.544
R 0.064 0.047 0.058 0.559 0.025 0.243 0.048 0.047 0.025 0.047 0.110 0.005 0.299 0.004 0.421 0.302
F1 0.116 0.089 0.105 0.506 0.046 0.259 0.088 0.088 0.047 0.087 0.176 0.010 0.387 0.007 0.424 0.389
SMD P 0.194 0.151 0.162 0.091 0.112 0.042 0.176 0.185 0.197 0.201 0.706 0.043 0.107 0.454 0.117 0.095
R 0.305 0.145 0.255 0.257 0.022 0.201 0.181 0.173 0.124 0.126 0.007 0.005 0.304 0.020 0.359 0.332
F1 0.237 0.148 0.198 0.135 0.037 0.069 0.178 0.179 0.152 0.155 0.014 0.009 0.158 0.039 0.176 0.148
ASD P 0.309 0.237 0.230 0.116 0.076 0.069 0.348 0.261 0.301 0.207 0.222 0.077 0.063 0.455 0.088 0.100
R 0.225 0.178 0.253 0.183 0.022 0.043 0.299 0.147 0.135 0.177 0.101 0.010 0.160 0.063 0.265 0.242
F1 0.219 0.158 0.184 0.107 0.031 0.047 0.258 0.177 0.163 0.146 0.129 0.017 0.085 0.108 0.125 0.131
Contextual4.9 P 0.584 0.067 0.555 0.044 0.051 0.049 0.101 0.264 0.446 0.119 0.670 0.041 0.234 0.577 0.072 0.029
R 0.733 0.340 0.696 0.235 0.356 0.364 0.559 0.713 0.656 0.656 0.591 0.069 0.559 0.393 0.174 0.053
F1 0.650 0.112 0.618 0.075 0.089 0.086 0.172 0.385 0.531 0.201 0.628 0.051 0.330 0.467 0.102 0.037
Contextual7.2 P 0.744 0.095 0.548 0.085 0.080 0.067 0.157 0.352 0.373 0.160 0.775 0.084 0.306 0.633 0.103 0.053
[…]